Circular No. 6703

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

              URL http://cfa-www.harvard.edu/cfa/ps/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

V723 CASSIOPEIAE T. Iijima and L. Rosino, Asiago Astrophysical Observatory, report:

“A spectrogram (range 460–710 nm, resolution 0.06 nm) of V723 Cas (N Cas 1995) was taken at Asiago Observatory on July 1.09 UT using the 1.82-m reflector (+ échelle spectrograph). Prominent emission lines of [Fe VII] at 572.1 nm (equivalent width −2.76 nm) and at 608.6 nm (EW = −7.04 nm) were detected in the spectrum. These lines were not seen in a spectrum taken on May 30.1 with the same instruments. The intensity of the emission line He II 468.6-nm relative to Hβ increased from 0.16 to 0.54 during the same period, while [O III] 500.7-nm was nearly stable and its intensity was ∼ 0.15 of Hβ in both spectra. Prominent emission lines in the last spectrum, in order of decreasing intensity, are as follows (line positions given in nm): Hα, Hβ, He II 468.6, He I 587.6, He I 706.5, [N II] 575.5, [Fe VII] 608.6, [O III] 500.7, He I 667.8, [Fe VII] 572.1, Si II 504.1, He II 541.2, [Fe VI] 517.7, [Fe II] 464.0, [Fe III] 527.0, He I 501.6, [S III] 631.0, [Ca V] 530.9, [Fe II] 497.3, [O III] 495.5, Si II 505.6, Si II 637.1, and [Fe VII] 515.8. About two years after the explosion, this slowest classical nova seems to have entered the nebular stage. The high intensities of Si II lines, however, are fairly unusual for spectra of classical novae in this stage.”

BL LACERTAE R. Hartman, D. Bertsch, S. Bloom, P. Sreekumar, and D. Thompson,

Goddard Space Flight Center, report: “The EGRET instrument on the Compton Gamma Ray Observatory detects a significant increase in BL Lac at γ-ray energies above 100 MeV, in observations made during a target of opportunity following the announcement of an increase of over 1 mag in the optical emission of this source (IAUC 6693, 6700). The average γ-ray flux above 100 MeV is 1.3 × 10^{−6} photons cm^{−2} s^{−1}, based on data for 2.3 days. The significance level of the detections is 6.6σ. The current flux level is 3.5 times that of the only other weak detection in January 1995. Prior to 1995, EGRET only detected upper limits from BL Lac.”

F. Ma and D. Barry, University of Texas, report that UBVRI images of the field of BL Lac were made with the McDonald Observatory 0.76m telescope between July 19.21 and 21.46 UT. The preliminary lightcurve shows that it might have peaked on July 19.220 with R = 12.62 ± 0.05. Between July 19.454 and 19.467, it faded ∼ 0.3 mag, from R = 13.08 ± 0.05 to 13.37 ± 0.05, in 19 min. Additional R magnitudes: July 20.2243, 13.2; 20.4694, 12.86; 21.238, 13.69; 21.455, 13.07.

1997 July 21  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1997 CBAT  Daniel W. E. Green